OpenStudy (anonymous):

(square root of -7) (square root of -28)

OpenStudy (igreen):

\(\sqrt{-7} \times \sqrt{-28} = -14\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @jh32991

OpenStudy (anonymous):

Thank you igreen!

OpenStudy (igreen):

No problem!